### v2

@@ 3.1 AdS case
$\ln Z_{n}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right)+\frac{4r^{2}}{\lambda_{\rm AdS}}\left(-1+\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}\right).$ (36)
@@
$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\log Z_{n}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right).$ (37)
The entanglement entropy between two antipodal points on a sphere reflects the number of degrees of freedom in the field theory defined on the sphere Casini:2012ei . Casini and Huerta, using the principles of unitarity and locality in quantum field theory (QFT), demonstrated that the $C$-function, which measures the degrees of freedom, exhibits a monotonically decreasing behavior along the renormalization group flow for unitary and Lorentz-invariant QFTs Casini:2004bw ; Barnes:2004jj ; Gukov:2015qea . This validates Zamolodchikov’s $C$-theorem Zamolodchikov:1986gt . Since the $T\bar{T}$ trace flow equation is equivalent to the holographic renormalization group flow equation, the renormalization group flow can be controlled by adjusting the radius of the sphere deBoer:1999tgo ; Verlinde:1999xm ; Heemskerk:2010hk ; Nishioka:2018khk ; Grieninger:2023knz ; Kamei:2024iei . In the AdS case, the $C$-function is given by
$C(r)=3rS_{A}^{\prime}(r)=c_{\rm AdS}\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\,.$ (38)
As illustrated in figure 3, the $C$-function decreases monotonically as $r$ decreases, consistent with the number of degrees of freedom decreases along the renormalization group flow. Additionally, as $r$ approaches infinity, the $C$-function converges to the central charge of the two-dimensional CFT, indicating that the renormalization group flow approaches the UV fixed point Nakayama:2013is . This behavior is in agreement with the $C$-theorem.
@@
The $C$-theorem can be proven using locality and unitarity, but the converse is not true. Therefore, the locality of the field theory is next examined through the strong subadditivity inequality and the boosted strong subadditivity inequality. The strong subadditivity inequality is given by
$S(A)+S(B)-S(A\cup B)-S(A\cap B)\geq 0,$ (39)
where the subsystems $A$ and $B$ are shown in figure 4(a). In the limit where $A\rightarrow B$ and $|A|=|B|=L$, the strong subadditivity inequality can be expressed in its infinitesimal form as
$L^{2}S^{\prime\prime}(L)\leq 0,$ (40)
where $L$ represents the size of the subsystems.
(a) strong subadditivity
(b) boosted strong subadditivity
Figure 4: The left figure shows two spatial regions $A$ and $B$ at $t=0$. The right figure considers the regions $A$ and $B$ after a boost, resulting in two spacelike regions.
A stronger criterion for proving locality, compared to the strong subadditivity, is the boosted strong subadditivity, which can be derived by applying a boost transformation to subsystems $A$ and $B$, as shown in figure 4(b). The infinitesimal form of the boosted strong subadditivity inequality is expressed as
$L^{2}S^{\prime\prime}(L)+LS^{\prime}(L)\leq 0.$ (41)
Since the subsystems correspond to great arcs between antipodal points on a sphere, their size is proportional to the sphere’s radius, i.e., $L=\pi r$. Consequently, the derivatives with respect to $L$ in the strong subadditivity and the boosted strong subadditivity inequalities are equivalent to derivatives with respect to $r$. Taking the second derivative of eq. (37) with respect to $r$ yields the strong subadditivity inequality
$r^{2}\partial^{2}_{r}S_{A}(r)=-\frac{c_{\rm AdS}}{3\left(1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}\right)^{\frac{3}{2}}}<0.$ (42)
This result shows that the strong subadditivity inequality is always satisfied in the AdS case, while the boosted strong subadditivity inequality is violated. This violation is consistent within the framework of $T\bar{T}$-deformed theories, where the boundary field theory resides on a finitely truncated surface. In such scenarios, the entanglement wedge of the boosted subsystem $A$ is not fully contained within the entanglement wedge of $A\cup B$ Lewkowycz:2019xse .
3.2 dS case
In the context of the dS spacetime, there are two primary holographic models: dS/CFT and dS/dS. The focus here is on the dS/dS correspondence in Euclidean signature, where the holographic boundary is a 2-dimensional sphere. The metric of the dS spacetime in global coordinates is given by
@@
where the radius of the boundary sphere is $r=\ell_{\rm dS}\cosh t$, which exceeds $\ell_{\rm dS}$ for any finite cutoff at $t=t_{c}$. In dS/dS coordinates, the metric of the dS spacetime is
$ds^{2}=\ell_{\rm dS}^{2}\left[dw^{2}+\sin^{2}w\left(-d\tau^{2}+\cosh^{2}\tau\,d\phi^{2}\right)\right],$ (44)
where the effective sphere radius is $r=\ell_{\rm dS}\sin w<\ell_{\rm dS}$. The point $r=\ell_{\rm dS}$ corresponds to the cosmological horizon of the dS spacetime. When $r>\ell_{\rm dS}$, the hypersurface $r=r_{c}$ is spacelike, which aligns with the standard dS/CFT duality. Conversely, when $r<\ell_{\rm dS}$, the hypersurface $r=r_{c}$ becomes timelike, corresponding to the dS/dS correspondence, as illustrated in figure 5.
Figure 5: The Penrose diagram for the dS spacetime. The orange solid line represents future infinity, while the orange dashed line represents the finite cutoff at $t=t_{c}$. The red squares represent the boundaries at $w=0,\pi$ in dS/dS coordinates. The blue solid line represents $w=\pi/2$, and the blue dashed line represents the finite cutoff at $w=w_{c}$.
@@
$\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{i4r^{2}}{\lambda_{\rm dS}}\left(-1+\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}\right).$ (46)
@@
$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right).$ (47)
In this case, the argument of the inverse hyperbolic function is greater than one, making $\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)$ a complex value. The positive branch of the inverse hyperbolic function is chosen to match the real part of the entanglement entropy in the dS/CFT. The final expression for the entanglement entropy in the dS/CFT correspondence is
$S_{A}=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$ (48)
where the real part agrees with the results in refs. Hikida:2022ltr ; Doi:2023zaf ; Chen:2023eic . As discussed in ref. Doi:2022iyj , the entanglement entropy in the dS/CFT can be interpreted as the pseudo entropy, which is viewed as the origin of the non-unitarity of the boundary field theory in the dS/CFT. Additionally, we still want to use the $C$-function to measure the degree of freedom of the field theory. However, due to the non-unitarity of the dual field theory, the $C$-function is also a purely imaginary number.
$C(r)=3rS_{A}^{\prime}(r)=-ic_{\rm dS}\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\,.$ (49)
For the undeformed CFT, we have $C(r\to\infty)=-ic_{\rm dS}=c$. But when $r\to\ell_{\rm dS}$, the imaginary part of the $C$-function diverges, as shown in figure 6.
@@
Figure 6: The $C$-function of the dS spacetime at $r>\ell_{\rm dS}$, and $\lambda=\frac{\lambda_{\rm dS}}{r^{2}}$. The red dashed line represents $r=\ell_{\text{dS}}$.
In the dS/dS case, the entanglement entropy and $C$-function have been analyzed in refs. Gorbenko:2018oov ; Dong:2018cuv , with the results given as
$S_{A}=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}\right),\quad C(r)=3rS_{A}^{\prime}(r)=c_{\rm dS}\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}.$ (50)
It is noted that when the effective radius is expressed as $r=\ell_{\rm dS}\sin w<\ell_{\rm dS}$, these results can be directly obtained through analytic continuation from the corresponding entanglement entropy and $C$-function in the dS/CFT correspondence. The $C$-function does decrease monotonically as $r$ decreases, and it still diverges at $r=\ell_{\rm dS}$, as shown in figure 7.
@@
Figure 7: The $C$-function of the dS spacetime at $r<\ell_{\rm dS}$.
By calculating the second derivative of the entanglement entropy with respect to the sphere radius $r$, it can be found that when the dual field theory resides on a timelike boundary, the entanglement entropy violates the strong subadditivity inequality,
$r^{2}\partial^{2}_{r}S_{A}(r)=\frac{1}{\left(\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1\right)^{\frac{3}{2}}}>0.$ (51)
Therefore, it implies that the dual field theory is non-local Kawamoto:2023nki ; Geng:2019ruz when the effective radius of the sphere is smaller than the cosmological horizon.
@@ 4 Entanglement entropy for arbitrary single spatial intervals
In section 3, the replica method is employed to calculate the entanglement entropy for a subsystem between two antipodal points on a sphere under $T\bar{T}$ deformation. This method applies only to subsystems along semicircular arcs between antipodal points. To analyze the entanglement entropy for subsystems in arbitrary single spatial intervals on a two-dimensional plane, ref. Deng:2023pjs introduced the CHM mapping, which transforms the subsystem into a semicircular arc on a two-dimensional sphere. The entanglement entropy can then be determined using the results obtained for the spherical case. For subsystems on more general manifolds, they can first be mapped onto the two-dimensional plane, after which the method from ref. Deng:2023pjs can be applied. The replica method is generalized in the following section to calculate the entanglement entropy for arbitrary single spatial intervals on the two-dimensional plane with $T\bar{T}$ deformation and extended to general manifolds, including cylinders and spheres.
@@ 4.1 The generalization of the replica method
The goal is to calculate the entanglement entropy of the $T\bar{T}$-deformed quantum field theory in a spatial subregion $A$ on an arbitrary two-dimensional manifold $\mathcal{M}$, where $A$ is defined by $t=0$ and $x\in[x_{a},x_{b}]$ in a chosen coordinate system. When $\mathcal{M}$ is a two-dimensional sphere and $A$ is specified by $\phi=\text{constant}$ and $\theta\in(0,\pi)$, the entanglement entropy can be computed using eqs. (29) and (30).
The entanglement entropy of $A$ can be obtained through the $C$-function. Differentiating $S_{A}$ with respect to $r$, we have
$=2\pi r^{2}\lim_{n\rightarrow 1}n^{2}\int_{0}^{\pi}\sin\theta d\theta\,\partial_{n}\text{tr}\,T,$ (52)
where, in the last two steps, it is assumed that the metric (27) is independent of the azimuthal angle $\phi$, and thus $T_{ab}$ is also independent of $\phi$. Solving the trace flow and conservation equations without assuming spherical symmetry must determine the energy-momentum tensor $T_{ab}$. For the $n$-replica sphere (27), the solution for the energy-momentum tensor is known Lewkowycz:2019xse and is given by
For a general manifold $\mathcal{M}$, we first map $\mathcal{M}$ to $\mathbb{S}^{2}$, ensuring that the image of the subsystem $A$ under this mapping is a meridian on $\mathbb{S}^{2}$. The mapped metric is conformally equivalent to the sphere, denoted as $\Omega\mathbb{S}^{2}$. After applying the $n$-replica trick, we assume the metric takes the form
where the conformal factor $\Omega(\theta,\phi)$ is constant at the branch points $\theta=0$ and $\theta=\pi$, denoted as $\Omega_{0}$ and $\Omega_{\pi}$, respectively. For this metric, eqs. (29) and (30) remain valid
$\frac{d}{dr}\ln Z_{n}$ $=-\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\frac{\delta\mathcal{L}}{\delta g_{ab}}\frac{dg_{ab}}{dr}=-\frac{1}{r}\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\sqrt{g}T^{ab}g_{ab}.$ (57)
Thus, we have
$r\frac{d}{dr}S(r)$ $=r\frac{d}{dr}\lim_{n\rightarrow 1}(1-n\partial_{n})\ln Z_{n}=\lim_{n\rightarrow 1}(1-n\partial_{n})r\frac{d}{dr}\ln Z_{n}$
$=-\lim_{n\rightarrow 1}(1-n\partial_{n})nr^{2}\int_{\Omega\mathbb{S}_{n}^{2}}\Omega^{2}(\theta,\phi)\sin\theta\,d\theta d\phi\,\text{tr}\,T$
Since the conformal factor $\Omega(\theta,\phi)$ depends on the azimuthal angle $\phi$, $T_{ab}$ cannot be assumed to depend solely on $\theta$. Therefore, the $\phi$ coordinate is not integrated in the final two steps. In standard CFT calculations, a conformal transformation is often applied to map $\Omega\mathbb{S}^{2}$ to the spherical metric, and conformal symmetry ensures that the results for $\Omega\mathbb{S}^{2}$ and $\mathbb{S}^{2}$ differ only by a function of the conformal factor. However, for the $T\bar{T}$-deformed field theory, conformal symmetry is absent, and the relationship between quantities before and after the conformal transformation is not determined. As discussed in ref. Lewkowycz:2019xse , for the partition function on the sphere (52), in the limit $n\rightarrow 1$, the main contribution to the integral (52) comes from a small neighborhood near the branch points, with radius $\epsilon\sim\sqrt{n-1}$. This property is expected to hold for any manifold $\mathcal{M}$, so the integral (58) is also performed only in small neighborhoods near the branch points. In these regions, the conformal factor $\Omega$ is constant, and the metric (56) can be approximated by the $n$-replica sphere with radius $r\Omega$. As a result, the conformal transformation can be regarded as an identity transformation. Therefore, partition functions on any manifold $\mathcal{M}$ can be calculated using the results on $\mathbb{S}^{2}$ for non-conformal field theories. Finally, we can get
$rS^{\prime}(r)$ $=-\lim_{n\rightarrow 1}(1-n\partial_{n})\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\sqrt{g}\,\text{tr}\,T$
$=\lim_{n\rightarrow 1}2\pi(r\Omega)^{2}\left(\int_{0}^{\theta_{0}}+\int_{\pi-\theta_{0}}^{\pi}\right)\sin\theta\,d\theta\,\partial_{n}\text{tr}\,T,$ (59)
where $\text{tr}\,T$ corresponds to the solution on the sphere with radius $r\Omega$.
When Poincaré coordinates are chosen in EAdS3, the holographic boundary corresponds to a two-dimensional complex plane with the metric
$ds_{2}^{2}=\frac{\ell_{\rm AdS}^{2}}{z_{c}^{2}}(dT^{2}+dX^{2}).$ (60)
The subregion $A$ is defined by $T=0$ and $X\in(-X_{a},X_{a})$. To compute the entanglement entropy for $A$, the complex plane is mapped onto a two-dimensional sphere such that the image of $A$ corresponds to the meridian on the sphere at $\phi=0$, as illustrated in figure 8. This mapping, known as the CHM mapping, is given by
which is conformally equivalent to a sphere with an effective radius $r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$, and the conformal factor $\Omega(\theta,\phi)=(1+\cos\phi\sin\theta)^{-1}$.
@@
Figure 8: Mapping from the plane to the sphere. The north and south poles of the sphere correspond to the endpoints of the interval $A$ on the complex plane. The subregion $A$ corresponds to the meridian at $\phi=0$, while the locally time-evolved subregions (shown as blue or yellow lines) correspond to other meridians on the sphere.
At this stage, the energy-momentum tensor for the metric (62) must be determined using the equations in (31), followed by evaluating the integral in (52). As previously noted, when computing the partition function on a general manifold $\mathcal{M}$, the primary contribution to the integral (52) comes from an infinitesimal neighborhood near the branch points. In these regions, the conformal factor $\Omega(\theta,\phi)$ can be approximated as 1, implying that the metric (62) effectively describes a sphere with radius $r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$. Thus, within the infinitesimal neighborhoods around the endpoints, transforming (60) to the sphere involves only a conformal transformation with a conformal factor of 1, effectively an identity transformation. Consequently, the partition function remains the same as that of the sphere, with the radius $r$ replaced by the effective radius $r_{\mathrm{eff}}$.
For the AdS/CFT, the deformation parameter $\lambda$ and the central charge $c$ in eq. (31) are $c_{\rm AdS}=\frac{3\ell_{\rm AdS}}{2G}$ and $\lambda_{\rm AdS}=8G\ell_{\rm AdS}$, respectively. Therefore, the solution for the energy-momentum tensor on the $n$-replica sphere is
${T^{\phi}}_{\phi}$ $=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\frac{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}}{\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}}\right)\,,$ (63)
where the effective radius is $r_{\rm eff}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$. Substituting this into eq. (52) and performing the integration near the endpoints yields the $C$-function as
Using the $C$-function, one can get the entanglement entropy of the subregion A as
$S_{A}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\right)\simeq\frac{c_{\mathrm{AdS}}}{3}\log\frac{2X_{a}}{z_{c}}+\frac{c_{\mathrm{AdS}}}{12}\frac{z_{c}^{2}}{X_{a}^{2}}.$ (65)
In this expression, the first term represents the entanglement entropy of the undeformed field theory in the interval $(-X_{a},X_{a})$ on the plane. The second term accounts for the correction introduced by the $T\bar{T}$ deformation. This result is consistent with the findings reported in ref. He:2023xnb .
For the dS/CFT, the deformation parameter $\lambda$ and the central charge $c$ in eq. (31) are given by $c\equiv-ic_{\rm dS}=-i\frac{3\ell_{\rm dS}}{2G}$ and $\lambda\equiv-i\lambda_{\rm dS}=-i8G\ell_{\rm dS}$, respectively. The solution for the energy-momentum tensor on the $n$-replica sphere
@@
where the effective radius is $r_{\rm eff}=\frac{\ell_{\rm dS}X_{a}}{\eta_{c}}$. Substituting this into eq. (52), we obtain the $C$-function is
$C(r)=-\frac{ic_{\rm dS}}{3}\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\,.$ (67)
The entanglement entropy is then given by
$S_{A}=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}\simeq-\frac{ic_{\rm dS}}{3}\log\frac{2X_{a}}{\eta_{c}}+\frac{ic_{\rm dS}\eta_{c}^{2}}{12X_{a}^{2}}+\frac{\pi c_{\rm dS}}{6}\,.$ (68)
In Appendix A.1, this quantity is computed using holographic methods by evaluating the area of the RT surface with a finite cutoff $\eta_{c}$. The result is consistent with eq. (68).
To generalize the method for computing subsystems’ entanglement entropy on arbitrary holographic boundaries, we note that any two-dimensional manifold is locally conformally equivalent to the plane. Thus, its line element can be expressed as

### v3

@@ 3.1 AdS spacetime
$\ln Z_{n}=\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right)+\frac{4r^{2}}{\lambda_{\rm AdS}}\left(-1+\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}\right).$ (37)
@@
$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\log Z_{n}=\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right).$ (38)
The entanglement entropy between two antipodal points on a sphere reflects the number of degrees of freedom in the field theory defined on the sphere Casini:2012ei . Casini and Huerta, using the principles of unitarity and locality in quantum field theory (QFT), demonstrated that the $C$-function, which measures the degrees of freedom, exhibits a monotonically decreasing behavior along the renormalization group flow for unitary and Lorentz-invariant QFTs Casini:2004bw ; Barnes:2004jj ; Gukov:2015qea . This validates Zamolodchikov’s $C$-theorem Zamolodchikov:1986gt . Since the $T\bar{T}$ trace flow equation is equivalent to the holographic renormalization group flow equation, the renormalization group flow can be controlled by adjusting the radius of the sphere deBoer:1999tgo ; Verlinde:1999xm ; Heemskerk:2010hk ; Nishioka:2018khk ; Grieninger:2023knz ; Kamei:2024iei . In the AdS case, the $C$-function can be constructed through the entanglement entropy.
$C(r)\equiv 3rS_{A}^{\prime}(r)=c_{\rm AdS}\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\,.$ (39)
As illustrated in figure 3, the $C$-function decreases monotonically as $r$ decreases. Additionally, as $r$ approaches infinity, the $C$-function converges to the central charge of the two-dimensional CFT, indicating that the renormalization group flow approaches the UV fixed point Nakayama:2013is . This behavior is in agreement with the $C$-theorem.
@@
3.2 dS spacetime
In the context of dS spacetime, two primary holographic models are considered: dS/CFT and dS/dS holography, as illustrated in figures 4(a) and 4(b). The dS/CFT duality is conventionally formulated in global dS coordinates, whose metric is expressed as
@@
where the $T\bar{T}$-deformed CFT resides on a boundary at a finite cutoff $t=t_{c}$ and the radius of the boundary sphere, $r=\ell_{\rm dS}\cosh t_{c}$, exceeds $\ell_{\rm dS}$ for any finite cutoff except $t_{c}=0$. For dS/dS holography, the metric in dS/dS coordinates is
$ds^{2}=\ell_{\rm dS}^{2}\left[dw^{2}+\sin^{2}w\left(-dt^{2}+\cosh^{2}t\,d\phi^{2}\right)\right].$ (41)
In the Euclidean signature, the dual field theory is defined on a two-dimensional sphere with a radius $r=\ell_{\rm dS}\sin w_{c}$. Notably, for any finite radial cutoff $w_{c}$, the sphere’s radius $r$ remains strictly less than $\ell_{\rm dS}$. The critical condition $r=\ell_{\rm dS}$ corresponds to the central slices of dS spacetime, as depicted by the orange and blue dashed lines in Figure 4. According to the surface/state correspondence Miyaji:2015yva , since the central slice in dS spacetime is an extremal surface devoid of internal entanglement, the dual field theory on it should exhibit non-locality, i.e., we expect the $C$-function to diverge.
(a) dS/CFT
(b) dS/dS
Figure 4: The Penrose diagram for de Sitter spacetime. 4(a)The orange solid line represents the finite cutoff at $t=t_{c}$, while the orange dashed line represents $t=0$. 4(b)The green squares represent the boundaries at $w=0,\pi$ in dS/dS coordinates. The blue dashed line represents $w=\pi/2$, and the blue solid line represents the finite cutoff at $w=w_{c}$.
dS/CFT:
@@
$\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{i4r^{2}}{\lambda_{\rm dS}}\left(-1+\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}\right).$ (43)
@@
$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right).$ (44)
In this case, the argument of the inverse hyperbolic function is greater than one, making $\tanh^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)$ a complex value. The positive branch of the inverse hyperbolic function is chosen to match the real part of the entanglement entropy in the dS/CFT. The final expression for the entanglement entropy in the dS/CFT correspondence is
$S_{A}=-\frac{ic_{\rm dS}}{3}\coth^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$ (45)
where the real part agrees with the results in refs. Hikida:2022ltr ; Doi:2023zaf ; Narayan:2022afv ; Narayan:2023zen . As discussed in ref. Doi:2022iyj , the entanglement entropy in the dS/CFT can be interpreted as the pseudo entropy, which is viewed as the origin of the non-unitarity of the boundary field theory in the dS/CFT. Additionally, we still want to use the $C$-function to measure the degree of freedom of the field theory. However, due to the non-unitarity of the dual field theory, the $C$-function is also a purely imaginary number.
$C(r)=-ic_{\rm dS}\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\,.$ (46)
For the undeformed CFT, we have $C(r\to\infty)=-ic_{\rm dS}=c$. But as $r\to\ell_{\rm dS}$, the imaginary part of the $C$-function diverges, as illustrated in figure 5, as anticipated.
@@
Figure 5: The $C$-function of the dS spacetime at $r>\ell_{\rm dS}$, and $\lambda=\frac{\lambda_{\rm dS}}{r^{2}}$. The red dashed line represents $r=\ell_{\text{dS}}$, that is $t_{c}=0$.
dS/dS:
In the dS/dS holography, the entanglement entropy and $C$-function have been analyzed in refs. Gorbenko:2018oov , with the results given as
$S_{A}=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}\right),\quad C(r)=c_{\rm dS}\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}.$ (47)
These two results can also be directly derived from expressions (44) and (46) within the dS/CFT duality by recognizing that the sphere’s radius is smaller than the de Sitter radius $\ell_{\rm dS}$. As anticipated, the $C$-function diverges at the ultraviolet boundary $w_{c}=\frac{\pi}{2}$, as illustrated in figure 6.
@@
Figure 6: The $C$-function of the dS spacetime at $r<\ell_{\rm dS}$. The red dashed line indicates the ultraviolet boundary at $w_{c}=\frac{\pi}{2}$.
Half-dS holography:
Recently, a novel holographic framework, half-dS holography, has been proposed Kawamoto:2023nki . This model introduces a timelike boundary into global dS spacetime. The induced metric on this boundary is given by
$ds^{2}=\ell_{\rm dS}^{2}(-dt^{2}+\cosh^{2}t\sin^{2}\theta_{0}d\phi^{2}).$ (48)
The dual field theory is defined on this timelike boundary, characterized by the condition $\theta=\theta_{0}$. To ensure that spacelike geodesics can probe deep into the bulk spacetime, the bulk region is restricted to $0<\theta<\theta_{0}$, as indicated by the shaded area in figure 7. Reference Kawamoto:2023nki proposes that the dual field theory in this model also exhibits non-locality. Specifically, when the timelike boundary is located at $\theta_{0}=\frac{\pi}{2}$, the subadditivity is consistently violated for $t\neq 0$, except for the subadditivity saturation at $t=0$. This position coincides with the ultraviolet boundary position $w_{c}=\frac{\pi}{2}$ in dS/dS holography, which corresponds to the black dashed line in figure 7 and the blue dashed line in figure 4(b), respectively.
Figure 7: Half-dS holography. Here, we set $\cos T=\frac{1}{\cosh t}$. The dual field theory is defined on the timelike boundary $\theta=\theta_{0}$, and the bulk spacetime is represented by the shaded region.
The dual field theories in dS/dS holography and half-dS holography are both defined on timelike boundaries within dS spacetime and exhibit non-locality at the central slice. To this end, we will extend the replica method for calculating entanglement entropy between antipodal points on a sphere to compute the entanglement entropy between any spatial region and its complement on a general manifold in Section 4. Subsequently, we will apply the replica method to dS/dS holography and half-dS holography in Section 5 to discuss the similarities and differences between these two holographic models.
@@ 4 Entanglement entropy for arbitrary single spatial intervals
In section 3, the replica method is employed to calculate the entanglement entropy for a subsystem between two antipodal points on a sphere under $T\bar{T}$ deformation. This method applies only to subsystems along semicircular arcs between antipodal points. While the entanglement entropy of a single interval subjected to $T\bar{T}$ deformation on a planar manifold can be determined through CHM mapping and endpoint integration of the entanglement region Lewkowycz:2019xse ; Deng:2023pjs , a comprehensive calculation of the entanglement entropy for an arbitrary single interval under $T\bar{T}$ deformation on spherical and cylindrical manifolds remains absent [^4] In ref. Grieninger:2019zts , the authors proposed an entanglement entropy expression for a general interval, formatted analogously to the entanglement entropy expression between antipodal points on a spherical manifold, by introducing an effective radius. However, a rigorous proof of this proposition is lacking.. In this section, we will extend the replica method to general manifolds, including spherical and cylindrical manifolds, to compute the $T\bar{T}$ corrections to the entanglement entropy for any arbitrary single interval.
@@ 4.1 The generalization of the replica method
In ref. Lewkowycz:2019xse , it is proposed that in the $T\bar{T}$-deformed theory, all dependence of the entanglement entropy on $\lambda$ originates from the contributions of the endpoints of the entanglement region. For the reader’s convenience, we provide a concise review of the relevant derivation in Appendix A and apply it to a planar manifold via the CHM map, yielding the expression for the entanglement entropy of a single interval in AdS Poincaré spacetime. Drawing inspiration from this, we can employ analogous steps in dS Poincaré spacetime to derive the entanglement entropy expression for a single spatial interval in the $T\bar{T}$-deformed CFT. For dS Poincaré spacetime, the metric is given by
$ds^{2}=\frac{\ell_{\rm dS}^{2}}{\eta^{2}}(-d\eta^{2}+dX^{2}+dT^{2})\,.$ (49)
In this context, the $T\bar{T}$-deformed CFT resides on the two-dimensional complex plane $\eta=\eta_{c}$. The subregion $A$ is defined by $X\in[-X_{a},X_{a}]$ at $T=0$. It is mapped via the CHM transformation to a meridian $\Phi=0$ on a sphere with an effective radius $r_{\rm eff}=\frac{\ell_{\rm dS}X_{a}}{\eta_{c}}$, as illustrated in figure 8.
@@
Figure 8: Mapping from the plane to the sphere. The north and south poles of the sphere correspond to the endpoints of the interval $A$ on the complex plane. The subregion $A$ corresponds to the meridian at $\Phi=0$.
The solution for the energy-momentum tensor on the $n$-replica sphere is given by
@@
where the effective radius is $r_{\rm eff}=\frac{\ell_{\rm dS}X_{a}}{\eta_{c}}$. So we can obtain the $C$-function of the dS Poincaré spacetime
$C(r)=-\frac{ic_{\rm dS}}{3}\frac{1}{\sqrt{1-\frac{\eta_{c}^{2}}{X_{a}^{2}}}}\,.$ (51)
Finally, the entanglement entropy is given by
$S_{A}=-\frac{ic_{\rm dS}}{3}\coth^{-1}\left(\frac{1}{\sqrt{1-\frac{\eta_{c}^{2}}{X_{a}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}\simeq-\frac{ic_{\rm dS}}{3}\log\frac{2X_{a}}{\eta_{c}}+\frac{ic_{\rm dS}\eta_{c}^{2}}{12X_{a}^{2}}+\frac{\pi c_{\rm dS}}{6}\,.$ (52)
In Appendix B.1, we obtain consistent results by computing the area of the RT surface in the bulk spacetime with a finite cutoff $\eta_{c}$.
To generalize the replica method for arbitrary holographic boundaries, we note that any two-dimensional manifold is locally conformally equivalent to the plane. Thus, its line element can be expressed as